What is the magnitude of the electric field at a point midway between a −6.5μC and a +8.2μC charge 9.6cm apart? Assume no other charges are nearby.

Express your answer using two significant figures.
Answer #1

a = distance between the two charges = 9.6 cm = 0.096 m

r = distance of each charge from the midpoint = a/2 = 0.096/2 = 0.048 m

Q1 = magnitude of first charge = 6.5 x 10-6 C

Q2 = magnitude of second charge = 8.2 x 10-6 C

E1 = electric field by charge Q1

E2 = electric field by charge Q2

Electric field by charge Q1 at the midpoint is given as

E1 = k Q1 /r2 = (9 x 109) (6.5 x 10-6)/(0.048)2 = 2.5 x 107 N/C

Electric field by charge Q2 at the midpoint is given as

E2 = k Q2 /r2 = (9 x 109) (8.2 x 10-6)/(0.048)2 = 3.2 x 107 N/C

Net electric field at the midpoint is given as

E = E1 + E2 = (2.5 x 107) + (3.2 x 107)

E = 5.7 x 107 N/C